# GuideGen: A Text-Guided Framework for Full-torso Anatomy and CT Volume Generation

### v2

@@ Abstract
The recently emerging conditional diffusion models seem promising for mitigating the labor and expenses in building large 3D medical imaging datasets. However, previous studies on 3D CT generation have yet to fully capitalize on semantic and textual conditions, and they have primarily focused on specific organs characterized by a local structure and fixed contrast. In this work, we present GuideGen, a controllable framework that generates anatomical masks and corresponding CT volumes for the entire torso—from chest to pelvis—based on free-form text prompts. Our approach includes three core components: a text-conditional semantic synthesizer for creating realistic full-torso anatomies; a contrast-aware autoencoder for detailed, high-fidelity feature extraction across varying contrast levels; and a latent feature generator that ensures alignment between CT images, anatomical semantics and input prompts. To train and evaluate GuideGen, we compile a multi-modality cancer imaging dataset with paired CT and clinical descriptions from 12 public TCIA datasets and one private real-world dataset. Comprehensive evaluations across generation quality, cross-modality alignment, and data usability on multi-organ and tumor segmentation tasks demonstrate GuideGen’s superiority over existing CT generation methods.
@@ 1 Introduction
Figure 1: Overview of GuideGen. Given a medical prompt, GuideGen is able to synthesize anatomical structures that accurately reflect demographic or clinical details described in the prompt in three stages. In Stage-I, GuideGen synthesizes discrete semantic volumes that conforms to spatial features designated in the medical prompt; In Stage-II, GuideGen deploys a contrast-aware autoencoding module to accommodate large CT volumes while extracting detailed texture from all contrast levels; In Stage-III, GuideGen combines information of semantic latents derived from anatomies in Stage-I, image latents extracted in Stage-II and textual latents from the input medical prompt.
@@
The acquisition of a large quantity of medical images and their corresponding labels has always been critical for modern medical image analysis. However, due to data privacy issues and laborious annotation work, such datasets are often unavailable publicly, undermining the performance of subsequent tasks [14]. Recently trending conditional generative methods offer a promising solution to these problems: By eliminating privacy concerns and human interventions, they can generate vivid samples to an arbitrary scale, which provide significant boost to the performance of downstream applications like image segmentation and classification [31, 54, 29, 25].
Current conditional generative approaches can be categorized based on the nature of the conditions they incorporate. Semantic conditions, including organ and tumor maps, can be effectively harnessed to generate images that adhere to specific locality constraints [66, 65, 29, 24, 68]. Textual conditions like medical prompts, which are semantically rich and descriptive, can also be employed to enhance the diversity of generated images [7, 4, 9, 71, 30, 40]. However, despite their advancements, these methods fail to leverage the merits of both types of constraints, resulting in the semantic-guided models being not versatile enough to synthesize samples to a greater diversity and text-guided models not accurate enough to fully reflect exact spatial relationships among anatomical structures. Realizing this disadvantage, MedSyn [64] proposes to adopt an text-guided anatomy-aware generation pipeline that generates CT images not only conforms to anatomical masks, but also to a free-text textual guidance on a specific pulmonary disease. However, MedSyn requires paired semantics and texts as conditional input, and it only synthesizes a local region of the body (chest), which greatly limits its versatility and downstream usability. Moreover, contrary to the field of 2D medical image analysis, how to effectively incorporate textual knowledge and align it with semantic information remains largely underexplored in 3D CT synthesis [23]. This deficiency of targeted researches significantly hinders the formulation of an end-to-end, i.e. text to samples of masks and images, generation framework of medical CT dataset.
@@
To overcome these limitations, we propose a novel framework GuideGen that synthesizes full-torso anatomies and CT images based on a single medical free-text input, which are constructed by incorporating patient demographic information and clinical knowledge for a potential tumor site. Naturally, GuideGen can synthesize paired samples that provides for dataset construction on downstream multi-organ and tumor segmentation tasks. It is also more user-friendly as it necessitates only an medical free-text input, empowering researchers to synthesize full-torso CT dataset with minor effort. Our GuideGen consist of three individual stages, separately for text-guided semantics synthesis, contrast-aware CT autoencoding and latent-guided CT generation. In Stage-I, we offer a text-conditional semantic synthesizer based on a categorical diffusion model to generate exact discrete label indices, unlike current practices [24] that suffer from an ambiguity problem (See Sec. 3.1). It also translates the implicit information in medical prompts onto concrete semantic conditions to better guide the subsequent CT generation. In Stage-II, we train a contrast-aware autoencoder to extract comprehensive anatomical features from multiple contrast levels, avoiding the detail degradation by truncating intensities to a specific range that favors a region of interest as per many current practices [8, 68]. In Stage-III, we utilize a diffusion latent generator that operates on the combined information of semantic latents derived from anatomies in Stage-I, image latents extracted in Stage-II and textual latents capturing demographic or clinical features in the input medical prompt.
@@
We train and validate GuideGen on an assembly of 12 public tumor datasets and one in-house colorectal cancer dataset, which contain the multi-modal information of CT images and medical descriptions. To testify GuideGen’s downstream usability, we further include 2 multi-organ segmentation datasets and 3 tumor segmentation datasets. Our framework achieves state-of-the-art results in sample quality, conditional consistency as well as downstream usability on segmentation tasks across multiple datasets.
Our contributions can be concluded as follows:
  • We propose GuideGen, a viable and novel solution capable of performing text-based full-torso anatomy and CT generation.
  • We introduce a text-conditional semantic synthesizer that avoids ambiguity in current 3D anatomy generations, a contrast-aware module to accommodate high intensity variations in full-torso CT volumes and a latent-guidance module to better preserve detailed semantic guidance.
  • We conduct thorough experiments comparing the samples generated by GuideGen with 6 state-of-the-art generative frameworks from their quality, conditional consistency and downstream usability, on all of which GuideGen achieves better performances.
@@ 2 Related Works
Generative frameworks
such as Energy-Based Models [17, 21], Generative Adversarial Networks [15, 61, 73], Normalizing Flows [22, 36], Variational Autoencoders [41], and Diffusion Models [67, 32, 33] encouraged myriad researches on image generation due to their high authenticity and variability. Among these studies, the most advantageous feature that has emerged is arguably the capacity to steer the generation process via user-defined conditions [56]. This capability allows for a high degree of precision and customization in creating images, responding directly to the specific requirements set by the user, while providing a way for the generated images to be used for dataset augmentation purposes [19, 35]. In this work, we extends the ability of current diffusion models for a text-guided anatomy and CT synthesis, which mitigates data scarcity often experienced in the field of medical image analysis.
Medical CT Synthesis
on which previous endeavors undoubtedly provide some performance boost on their tasks of choice [25, 39, 13, 45, 29]. Despite their pioneering efforts, their work is confined to specific organs, which renders repeatedly retraining the network a necessity for each specific organ of interest. A lack of external guidance also underscores the need for a more flexible approach that conforms to not only descriptive texts but also to detailed semantics. Although recent work has begun to focus on these problems [64, 40], their generation processes require paired semantic and textual input to generate regional patches, which limits its downstream applicability. While a step forward, this indicates room for further development to achieve a more automated and generalized solution. This paper intends to provide a text-guided full-torso generative framework that aids to a wide range of downstream tasks.
@@ 3 Method
As illustrated in Fig. 1, GuideGen differs from current text-conditional medical image synthesizers [7, 23, 64] in that it disentangles the generative process into three stages separately for text-based anatomy placement, contrast-aware feature extraction and latent-guided image generation to address the multitude of labels, diversity of intensities and the incorporation of textual constraints in full-torso anatomy and CT volumes. We will elaborate on the specifics of our designs in the following sections.
3.1 Stage-I: Text-conditional Semantic Synthesizer
Current generative frameworks that uses dedicated semantic synthesizers [24, 10] unintentionally introduce ambiguities as they struggle to capture the sharp transitions in labels near semantic boundaries due to an inaccurate data modeling. They also fail to provide an interface for more fine-grained and versatile control via free-form texts. Contrary to these methods, our Text-Conditional Semantics Synthesizer (TCSS) provides a text-based ambiguity-free solution by building upon a categorical diffusion model [28, 70] which directly assumes a discrete formulation. As shown in Fig. 1(a), at its core, we train a categorical diffusion $p_{\theta}$ parameterized by $\theta$ that models the underlying distribution of mask volumes $\mathbf{m}$. At the first timestep, the diffusion variable $\mathbf{x}_{0}=\mathbf{m}\in\{1,\cdots,N\}^{H\times W\times D}$, where $H$, $W$, $D$ separately denotes mask height, width and depth. In the forward process, $\mathbf{x}_{0}$ is gradually transformed to a categorical noise $\mathbf{x}_{T}\sim\mathcal{C}_{N}(\mathbf{x}_{T};\mathbf{1}/N)$ in $T$ timesteps under a noise schedule $\beta_{1:T}$, where $\mathcal{C}_{N}$ denotes a categorical distribution of $N$ categories (semantic classes). The forward probability transition can be described by
@@ 3.1 Stage-I: Text-conditional Semantic Synthesizer
where the superscript $i$ is the voxel index and $\mathbf{e}(\cdot)$ is a function that returns one-hot probability vectors from a categorical input. For compactness and ease of reading, from now on we will consider the diffusion process on an image with only one voxel to lose the superscript $i$, i.e. $H=W=D=1$. Supposing voxels are independent, the formula derived below can be safely extended to images of any size. From the properties of Markov chains [28], the forward process of $p_{\theta}$ can be described as
@@
To extract relevant information in the medical prompt for semantic placement, we use a text encoder $\mathcal{E}_{T}$ pretrained on medical knowledge [62] to map the original prompt onto a latent space and perform cross-attention to guide anatomical generation by extracting spatial relation among anatomies. We interleave these cross-attention layers with a 3D U-Net [11] backbone to construct $p_{\theta}$. Specifically, for each intermediate layer $l$ in the 3D U-Net, the latent vector $\mathbf{z}_{l}$ is first computed by
$\mathcal{Q}=W^{q}(\mathbf{z}_{l-1}),\mathcal{K}=W^{k}(\mathcal{E}_{T}(\mathbf{p})),\mathcal{V}=W^{v}(\mathcal{E}_{T}(\mathbf{p}))$ (3)
@@
where $W^{q}$, $W^{k}$, and $W^{v}$ are learnable projection mappings, $\mathbf{p}$ is the input medical prompt and $d$ is the context dimension. We train $p_{\theta}:[0,1]^{N}\rightarrow[0,1]^{N}$ by minimizing the Kullback-Leiber divergence between posterior and generator distributions across the diffusion process as in [27]. By denoting the posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\frac{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})q(\mathbf{x}_{t-1}|\mathbf{x}_{0})}{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\in[0,1]^{N}$ as $\boldsymbol{\pi}_{t}(\mathbf{x}_{0})$, the overall training objective is
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})\right)\right],$ (5)
@@
To avoid training instability, we make a similar reparameterization as in DDPM [27] for enhanced performance by training a generator model on the same space as the input $\mathbf{x}_{0}$. Instead of training a model $p_{\theta}(\mathbf{x}_{t},\mathbf{p})$ that models on the distribution $\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})$, we will train and draw samples from its reparameterized version, denoted as $f_{\theta}=\mathcal{C}_{N}(\hat{\mathbf{x}}_{0};f_{\theta}(\mathbf{x}_{t},\mathbf{p}))$. The training loss $\mathcal{L}$ needs to be modified accordingly: From the total probability formula, denoting $\mathbf{A}=(\boldsymbol{\pi}_{t}(1),\cdots,\boldsymbol{\pi}_{t}(N))$, we have $p_{\theta}=\mathbf{A}f_{\theta}$. The voxel-wise loss for TCSS can be written as
=$\begin{split}\mathcal{L}&{}_{TCSS}=\\ &\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};\mathbf{A}f_{\theta})\right)\right].\end{split}$ (6)
Under real settings where an image actually contains multiple voxels, the total loss $\mathcal{L}_{TCSS}$ is averaged over each individual voxel.
At inference, TCSS samples from the distribution $f_{\theta}$ and computes the next denoising step by sampling on the generator distribution $S(\mathbf{A}f_{\theta}(\hat{\mathbf{x}}_{t},\mathbf{p}))$, where $S$ is the sampling scheme. Similar to [70], we sample by probability for timesteps $2\sim T$ and take the index with max probability on the final denoising step to add stability to the generation process.
3.2 Stage-II: Contrast-aware Autoencoder
Similar to [9, 23], we use a pair of image autoencoder to extract high-level texture details from CT volumes while lowering memory demands to incorporate larger CT volumes. As shown in Fig. 1(b), similar to physicians examining different anatomical structures using different intensity windows, our Contrast-aware Autoencoder (CAE) strives to accurately accommodate the entire dynamic range of CT intensities with the help of a contrast-aware rescaler $h$. Formally,

### v3

# GuideGen: A Text-Guided Framework for Paired Full-torso Anatomy and CT Volume Generation

@@ Abstract
The recently emerging conditional diffusion models seem promising for mitigating the labor and expenses in building large 3D medical imaging datasets. However, previous studies on 3D CT generation primarily focus on specific organs characterized by a local structure and fixed contrast and have yet to fully capitalize on the benefits of both semantic and textual conditions. In this paper, we present GuideGen, a controllable framework based on easily-acquired text prompts to generate anatomical masks and corresponding CT volumes for the entire torso—from chest to pelvis. Our approach includes three core components: a text-conditional semantic synthesizer for creating realistic full-torso anatomies; an anatomy-aware high-dynamic-range (HDR) autoencoder for high-fidelity feature extraction across varying intensity levels; and a latent feature generator that ensures alignment between CT images, anatomical semantics and input prompts. Combined, these components enable data synthesis for segmentation tasks from only textual instructions. To train and evaluate GuideGen, we compile a multi-modality cancer imaging dataset with paired CT and clinical descriptions from 12 public TCIA datasets and one private real-world dataset. Comprehensive evaluations across generation quality, cross-modality alignment, and data usability on multi-organ and tumor segmentation tasks demonstrate GuideGen’s superiority over existing CT generation methods. Relevant materials are available at https://github.com/OvO1111/GuideGen.
@@ 1 Introduction
Figure 1: Overview of GuideGen’s training pipeline: (a) Firstly, GuideGen learns to generate discrete semantic volumes that conforms to spatial features designated in the medical prompt (See Sec.3.1); (b) Secondly, GuideGen deploys a pyramidal autoencoding scheme to incorporate mask knowledge and reconstruct fine CT details with a high dynamic range (See Sec.3.2); (c) Finally, GuideGen combines the semantic latents derived from (a), image latents extracted in (b) and textual latents from the medical prompt to synthesize full-torso CT images (See Sec.3.3); (d) The internal structure of our knowledge injection module for extracting task-specific features from a structured input used in (a) and (c).
@@
The acquisition of a large quantity of medical images and their corresponding labels has always been critical for modern medical image analysis. However, due to data privacy issues and laborious annotation work, such datasets are often unavailable publicly, undermining the performance of subsequent tasks (dai2023efficient). Recently trending conditional generative methods offer a promising solution to these problems: By eliminating privacy concerns and human interventions, they can generate vivid samples to an arbitrary scale, which provide significant boost to the performance of downstream applications like image segmentation and classification (huang2024label; hashmi2024xreal).
Current conditional generative approaches can be categorized based on the nature of the conditions they incorporate. Semantic conditions, like organ and tumor maps, can be harnessed to generate images that adhere to specific locality constraints (yao2021label; yang2023diffusion; hu2023label; han2023medgen3d; guo2024maisi). Textual conditions, on the other hand, have the advantage of enhanced generation diversity and reduced human effort at sample synthesis thanks to the versatility of natural language (chambon2022roentgen; balaji2022ediffi; cho2024medisyn; huang2024chest; guo2025text2ct). However, current advancements that adopts conditional pipelines seldom leverage the merits from both categories. Consequently, semantic-guided models typically exhibit limited sample diversity and depend on comprehensive anatomy masks that are costly to curate. Conversely, text-guided models, though much more flexible, struggle to fully capture exact spatial relationships among anatomical structures. Realizing this disadvantage, Kim et al. (kim2024controllable) proposes to use a text-guided controllable generation pipeline that generates CT images not only conforms to a general semantics, but also to a textual guidance on image specifics. However, it requires a paired input of semantic and textual guidance as input, which greatly limits its applicability. MedSyn (xu2024medsyn) alleviates the need of paired input at inference by setting a null semantic input and recovering CT from a learned joint distribution. However, it remains unclear whether its chest-only pipeline can broadcast to the entire torso with more complex anatomical structures. Also, current text-guided generative pipelines largely focus on straightforward usages in classification tasks, and their potential for downstream segmentation remains underexplored, which is both critical in clinical treatment planning and accurate diagnosis (xu2024advances).
@@
To this end, we propose GuideGen, a novel framework that synthesizes full-torso anatomies and CT images based on medical strctured inputs. Naturally, GuideGen can synthesize paired samples that provides for dataset construction on downstream segmentation tasks. It is also more user-friendly than mask-based competitors as it necessitates only a textual input, empowering researchers to synthesize full-torso CT dataset with minor effort. GuideGen first offers a text-conditional semantic synthesizer based on a categorical diffusion model to generate discrete label indices, unlike current practices (han2023medgen3d) that suffer from an ambiguity problem (See Sec. 3.1). We also develop a knowledge injection module to help translate implicit information in medical prompts. Secondly, we train an anatomy-guided autoencoder that extracts comprehensive anatomical features from multiple contrast levels, avoiding the detail degradation by truncating intensities to a specific range that favors a region of interest (chen2024towards; yu2024ct) while preserving small organ and tumors described in the semantic condition. Finally, we utilize a diffusion latent generator that operates on the combined space of generated semantics and input textual features to recover valid image latents.
@@
We train and validate GuideGen on an assembly of 12 public tumor datasets and one in-house colorectal cancer dataset, which contain the multi-modal information of CT images and medical descriptions. To testify GuideGen’s downstream usability for segmentation tasks, we further include 2 multi-organ segmentation datasets and 3 tumor segmentation datasets. Our framework achieves state-of-the-art results in sample quality, conditional consistency as well as downstream usability on segmentation tasks across multiple datasets, all with the help of single textual inputs.
@@ 2 Related Works
Generative frameworks, such as Energy-Based Models (du2019implicit; guo2023egc), Generative Adversarial Networks (de2021impact; zhou2023gan), Normalizing Flows (hajij2022normalizing; jeevan2024normalizing), Variational Autoencoders (kingma2022autoencoding), and Diffusion Models (yoon2023sadm; hung2023med; iglesias2024generation) encouraged myriad researches on image generation due to their high authenticity and variability. Among these studies, the most advantageous feature that has emerged is arguably the capacity to steer the generation process via user-defined conditions (po2024state). This capability allows for a high degree of precision and customization in creating images, responding directly to the specific requirements set by the user, while providing a way for the generated images to be used for dataset augmentation purposes (fang2024data; islam2024diffusemix). In this work, we extend the ability of current diffusion models for a text-guided anatomy and CT synthesis, which mitigates data scarcity experienced in the field of medical image analysis.
Conditional Medical CT Synthesis, on which previous studies delivered convincing gains on selected downstream tasks (hashmi2024xreal; kazerouni2023diffusion; colleoni2024guided; konz2024anatomically; hu2023label; guo2024maisi). Despite their pioneering efforts, their work mostly rely on external semantic guidance at inference, which is not costly to attain at clinical practices. On the other hand, text-based generation pipelines (pinaya2022brain; hamamci2025generatect; guo2025text2ct) usually supplies too little information to yield high-quality volumes and can only benefit downstream classification tasks. Although recent work has begun to combine the merits from both worlds (kim2024controllable; xu2024medsyn), they require paired inputs at inference or generate only regional patches and cannot provide for downstream segmentation tasks. While a step forward, this indicates room for further development to achieve a more automated and generalized solution. This paper intends to provide a text-guided full-torso generative framework that aids to a wider range of downstream tasks.
@@ 3 Method
As illustrated in Fig. 1, GuideGen separates the generative process into three stages to generate full-torso anatomy and CTs for downstream tasks with text-only inputs. We will elaborate on our designs in the following sections.
3.1 Text-conditional Semantic Synthesizer
Ambiguity-reducing Categorical Modeling Current generative frameworks that uses dedicated semantic synthesizers (han2023medgen3d; chu2024anatomic) unintentionally introduce ambiguities as they struggle to capture the sharp transitions in labels near semantic boundaries due to an inaccurate data modeling. Contrary to these methods, our Text-Conditional Semantics Synthesizer (TCSS) provides a valid solution to reduce ambiguity by building upon a categorical diffusion model (hoogeboom2021argmax; zbinden2023stochastic) which directly assumes a discrete formulation. As shown in Fig. 1(a), at its core, we train a categorical diffusion $p_{\theta}$ parameterized by $\theta$ that models the underlying distribution of mask volumes $\mathbf{m}$. At the first timestep, the diffusion variable $\mathbf{x}_{0}=\mathbf{m}\in\{1,\cdots,N\}^{H\times W\times D}$, where $H$, $W$, $D$ separately denotes mask height, width and depth. In the forward process, $\mathbf{x}_{0}$ is gradually transformed to a categorical noise $\mathbf{x}_{T}\sim\mathcal{C}_{N}(\mathbf{x}_{T};\mathbf{1}/N)$ in $T$ timesteps under a noise schedule $\beta_{1:T}$, where $\mathcal{C}_{N}$ denotes a categorical distribution of $N$ categories (semantic classes). The forward probability transition can be described by
@@ 3.1 Text-conditional Semantic Synthesizer
where the superscript $i$ is the voxel index and $\mathbf{e}(\cdot)$ is a function that returns one-hot probability vectors from a categorical input. For compactness and ease of reading, from now on we will consider the diffusion process on an image with only one voxel to lose the superscript $i$, i.e. $H=W=D=1$. Supposing voxels are independent, the formula derived below can be safely extended to images of any size. From the properties of Markov chains (hoogeboom2021argmax), the forward process of $p_{\theta}$ can be described as
@@
Knowledge Injection To provide an interface for more fine-grained and versatile control via structured prompts, we extract relevant information in the medical prompt for semantic placement. Conventionally, it is widely adopted to use a dedicated text encoder $\mathcal{E}_{T}$ pretrained on medical knowledge (sun2021ernie) to map the original prompt $\mathbf{p}$ onto a latent space and perform cross-attention with the generation backbone (chambon2022roentgen; xu2024medsyn; hamamci2025generatect). However, this alone is often insufficient to steer the model’s focus to the most pertinent descriptions in input medical prompts. For instance, the model may fail to prioritize critical information like tumor location over less relevant descriptors, such as race and gender, during semantic synthesis. Recognizing this, we opt to use a separate module to extract task-specific features from encoded latents.
Specifically, as shown in Fig. 1(d), knowledge injection works by allowing learnable task-specific generation queries $Q$ to interact with a series of transformer decoder blocks to retrieve relevant responses $R_{\rm task}\in\mathbb{R}^{N\times C}$ for mask or image generative tasks from encoded medical prompts $\mathcal{E}_{T}(\mathbf{p})$. In addition, since different layers in the generative backbone perceive different levels of semantic information, we derive layer-wise responses $R_{\rm layer}\in\mathbb{R}^{N\times(C\times L)}$ focusing on either global anatomies or local structures from $R_{\rm task}$. $N$, $C$, $L$ separately denote the number of query tokens, latent dimension and network depth of $p_{\theta}$. This layer-wise guidance $R_{\rm layer}$ is then injected into the diffusion backbone $p_{\theta}$: For each intermediate layer $l$ in the diffusion backbone $p_{\theta}$, denoting its textual guidance as $R_{\rm layer}^{l}$, the generative latents $\mathbf{z}_{l}$ at layer $l$ is computed by
$\mathcal{Q}=W^{q}(\mathbf{z}_{l-1}),\mathcal{K}=W^{k}(R_{\rm layer}^{l}),\mathcal{V}=W^{v}(R_{\rm layer}^{l})$ (3)
@@
where $W^{q}$, $W^{k}$, and $W^{v}$ are learnable projections and $d$ is the context dimension.
Training Objective We train $p_{\theta}:[0,1]^{N}\rightarrow[0,1]^{N}$ by minimizing the Kullback-Leiber divergence between posterior and generator distributions across the diffusion process as in (ho2020denoising). Denoting the posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\frac{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})q(\mathbf{x}_{t-1}|\mathbf{x}_{0})}{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\in[0,1]^{N}$ as $\boldsymbol{\pi}_{t}(\mathbf{x}_{0})$, the overall training objective is
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{\rm KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})\right)\right],$ (5)
@@
To avoid training instability, we make a similar reparameterization as in DDPM (ho2020denoising) for enhanced performance by training a generator model on the same space as the input $\mathbf{x}_{0}$. Instead of training a model $p_{\theta}(\mathbf{x}_{t},\mathbf{p})$ that models on the distribution $\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})$, we will train and draw samples from its reparameterized version, denoted as $f_{\theta}=\mathcal{C}_{N}(\hat{\mathbf{x}}_{0};f_{\theta}(\mathbf{x}_{t},\mathbf{p}))$. The training loss $\mathcal{L}$ needs to be modified accordingly: From the total probability formula, denoting $\mathbf{A}=(\boldsymbol{\pi}_{t}(1),\cdots,\boldsymbol{\pi}_{t}(N))$, we have $p_{\theta}=\mathbf{A}f_{\theta}$. The voxel-wise loss for TCSS can be written as
$\mathcal{L}_{\rm 1}=\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{\rm KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};\mathbf{A}f_{\theta})\right)\right].$ (6)
Under real settings where an image actually contains multiple voxels, $\mathcal{L}_{\rm 1}$ is averaged over each individual voxel.
At inference, TCSS samples from the distribution $f_{\theta}$ and computes the next denoising step according to a sampling scheme $S$ on the generator distribution $S(\mathbf{A}f_{\theta}(\hat{\mathbf{x}}_{t},\mathbf{p}))$. Similar to (zbinden2023stochastic), we sample by probability for timesteps $2\sim T$ and take the index with max probability on the final denoising step to add stability to the generation process. Finally, generated masks can be retrieved via $\hat{\mathbf{m}}=\rm{argmax}(\hat{\mathbf{x}}_{0})$ along the channel dimension.
3.2 Anatomy-aware HDR Autoencoder
Anatomy Preservation Similar to previous works (cho2024medisyn; hamamci2025generatect), we use a pair of image autoencoder to extract high-level texture details from CT volumes while lowering memory demands to incorporate larger CT volumes. However, current practices that directly apply off-the-shelf autoencoder architectures could undermine the preservation of spatial relations among anatomies of different sizes. For example, a small tumor visible in a higher resolution input may not be as readily recognized by the decoder half or the subsequent diffusion model in the latent space of the autoencoder. Therefore, we propose to use semantic masks to constrain the autoencoding process and mitigate this undesirable effect. As illustrated in Fig. 1(b), we follow a pyramidal approach to resample the semantic information to the resolution of layer-wise latents in the autoencoder and concatenate them in each layer. This helps the model focus on semantic details at encoding and reconstruct semantically accurate images in the decoding process.
HDR Accommodation Secondly, unlike concurrent works that truncate input CT volumes to a limited intensity range (yu2024ct; chen2024towards; zhuang2023semantic), we aim to accommodate the entire dynamic range of intensities in full-torso CTs to incorporate details for each organ. Similar to physicians examining different anatomies using different intensity windows, we devise an intensity transformation module $h$ and constrain the reconstruction process under the full spectrum of intensity ranges. Formally,